Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-03-16
ISS On-Orbit Status 03/16/08

Sunday - 1J/A Flight Day 6/7 (FD6/7).

Ahead: Week 22 of Increment 16. Crew sleep cycle today: Sleep 7:00am -3:30pm; wake 3:30pm -6:00am tomorrow.

More good news!

SPDM Dextre was checked out and is working nominally with both arms. (The waist-up-only robot from Canada arrived in space in nine separate pieces that are being assembled in the current spacewalks. Each of the two arms has seven joints; in addition, SPDM (Special Purpose Dexterous Manipulator) Dextre can pivot at the 'waist'. Its grippers (hands) have built-in socket wrenches, cameras & lights. Only one arm is movable at a time, to keep the robot stable and avoid a two-arm collision. Dextre can be attached to MT (Mobile Transporter) to translate along the stations rail tracks, or alternately to the SSRMS (Space Station Remote Manipulator System) to swing to places where the railcart can't go. What a supercool helper!)

EVA-2 was completed successfully by Rick Linnehan and Mike Foreman in 7h 8m, accomplishing all its objectives.
(During the spacewalk, Linnehan (EV1) & Foreman (EV2) -

Removed SPDM Arms #1 & #2 from the SLP FSE (Spacelab Pallet Flight Support Equipment); the arms were then temporarily stowed on an SLP bracket. During removal of the EDFs (Expandable Diameter Fasteners) from both arms, the crew encountered difficulty in releasing the fasteners. A pry bar was required on both arms to finally release the fasteners,- a critically important and fortunate accomplishment;
Mated both Arm #1 & Arm #2 to Dextre's main body; and
Removed several of the 19 MLI (Multi-Layered Insulation) thermal covers from Arms #1 & #2; teams are still assessing what thermal covers are left and which covers need to be removed on EVA-3 tomorrow night.
Official start time of the spacewalk was 7:48pm EDT, about 35 min ahead of timeline, and it ended at 2:56am. Total EVA duration (PET = Phase Elapsed Time) was 7h 8min. It was the 106th spacewalk for ISS assembly & maintenance and the 78th from the station (56 from Quest, 22 from Pirs, 28 from Shuttle) totaling 479h 30min, the 10th for Expedition 16 (totaling 71h 38min) and the 7th so far this year. After today's EVA, a total of 133 spacewalkers (101 NASA astronauts, 21 Russians, and 11 astronauts representing Japan-1, Canada-4, France-1, Germany-2 and Sweden-3) have logged a total of 667h 52min outside the station on building, outfitting and servicing. It was also the 128th spacewalk involving U.S. astronauts.)

After the EVA, SSRMS re-grappled SPDM and power was re-applied. This allowed the SPDM arms and body to be brought to life (i.e., operational status) for the first time. Diagnostic tests were successfully performed overnight on both arms. Dextre's final assembly is scheduled to be completed tomorrow night on EVA-3. On the following day (FD9), SPDM will be grappled by the SSRMS and relocated to the U.S. Lab PDGF (Power & Data Grapple Fixture).

After the spacewalkers' ingress this morning at 2:56am, post-EVA activities by Peggy Whitson, Rick Linnehan, Mike Foreman and Bob Behnken and Leo Eyharts in the Airlock (A/L) consisted of recharging the EMU/spacesuits with water from PWR (Payload Water Reservoir), then reconnecting the LTAs (Lower Torso Assemblies) to the EMUs and capping the UIA (Umbilical Interface Assembly), initiating and monitoring the regeneration of the METOX (Metal Oxide) CO2 filter canisters, initiating the discharge/recharge process on the EMU batteries in the BSA (Battery Stowage Assembly), and reconfiguring EVA tools. (A temporary blower failure of the CDRA (Carbon Dioxide Removal System) at 6:35am had no impact on METOX regeneration. CDRA was reactivated at 7:08am and is currently running.)

During the EVA, CDR Whitson and MS3 Doi worked in the JAXA JLP (JEM Experiment Logistics Module Pressurized Section) to complete its outfitting and preparation for the arrival of the JEM (Japanese Experiment Module) 'Kibo' later this year. (Peggy and Takao rotated the module racks to obtain temperature readings of its shell (required to support an engineering evaluation of the thermal math models in case the JLP trunnion covers are not installed during one of the next EVAs), then released rack launch restraints, removed rack closeouts, installed pivot fittings and K-Bar assemblies and secured knee braces for the JEMRMS (JEM Robotic Manipulator System), ICS (Inter Orbit Communication System), Saibo, Ryutai and the JRSR (JEM Resupply Stowage Rack) racks. Rack translations within the JLP are planned during Flight 1J after the JLP is relocated to its final location on the JEM.
Background: Saibo ('living cell') is a Japanese multipurpose experiment/payload rack system on the ISS that transports, stores and supports subrack facilities such as the CB (Clean Bench) and CBEF (Cell Biology Experiment Facility) equipment by providing structural interfaces, power, data, cooling, water and other items needed to operate science experiments in microgravity. Ryutai ('fluid') is a Japanese multipurpose experiment/payload rack system to support the FPEF (Fluid Physics Experiment Facility), SCOF (Solution Crystallization Observation Facility), PCRF (Protein Crystallization Research Facility) and the IPU (Image Processing Unit) by providing structural interfaces, power, data, cooling, water and other items needed to operate science experiments in micro-G.)

In the Service Module (SM), FE-1 Malenchenko deinstalled the JAXA-3DPC (Japan Aerospace Exploration Agency-3D Photon Crystals) crystal growth experiment and transferred it to the Shuttle for return to Earth. (The experiment had been running since 1/22 and was turned off by Yuri on 1/31. Its purpose was to grow photonic crystals in microgravity using particles in electrolytic solutions that will be fixed using ultraviolet light in a process referred to as photocuring, after self-organization and ordering of colloid nanoparticles in an electrolyte solution with subsequent fixation in an elastic gel matrix.)

Malenchenko also unstowed a spare BRPK air/liquid condensate separator unit, flushed it out with clean water (to remove any sediments that may have accumulated as a result of length storage) and installed it in the BRPK-1 line of the SRV-K2M Condensate Water Recovery System.

As the spacewalk went on, Eyharts set up and conducted the periodic (monthly) O-OHA (On-Orbit Hearing Assessment) test, a 30-min NASA EHS (Environmental Health Systems) examination to assess the efficacy of acoustic countermeasures, using a special MEC (Medical Equipment Computer) laptop application. It was Leo's second onboard O-OHA session. (The O-OHA audiography test involves minimum audibility measurements for each ear over a wide range of frequencies (0.25-10 kHz) and sound pressure levels, with the crewmembers using individual-specific Prophonics earphones, Bose ANC headsets and the SLM (sound level meter). To conduct the testing, the experimenter is supported by special EarQ software on the MEC, featuring an up/down-arrow-operated slider for each test frequency that the crewmember moves to the lowest sound pressure level at which the tone can still be heard. The baseline test is required not later than about Flight Day 14 for each new Expedition and is then generally performed once per month. Note: There have been temporary hearing deficits documented on some U.S. and Russian crewmembers, all of which recovered to pre-mission levels.)

Later, Eyharts also installed 22 newly arrived white U.S. RAM (Radiation Area Monitor) dosimeters in the ISS, documenting their deployment photographically, and collected 18 used blue RAMs deployed on 8/13/07, for return on 1J/A.

Before sleeptime, Yuri successfully repaired the broken Multimeter instrument. (The activities began with FE-1 setting up the MWA (Maintenance Work Area), followed by the soldering of new battery terminal wires to the Multimeter.)

For both crews, currently asleep, workday begins today at 3:30pm EDT.

After crew wakeup, Leo will remove the deactivated InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions) hardware from the MSG (Microgravity Science Glovebox) for stowage.

Before breakfast, as suggested on her discretionary 'job jar' task list, Peggy Whitson will complete another 'overnight' run with the SLEEP experiment (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) software for data logging and filling in questionnaire entries in the experiment's laptop session file on the HRF-1 laptop for downlink. (To monitor the crewmember's sleep/wake patterns and light exposure, Peggy put on a special Actiwatch device which measures the light levels encountered by him as well as his patterns of sleep and activity throughout the Expedition. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew's discretionary 'job jar' task list.)

First thing for Yuri after wakeup will be to set up the hardware for the Russian MBI-21 PNEVMOKARD experiment and conduct the session, his fourth (which forbids moving or talking during data recording). The experiment is controlled from the RSE-Med A31p laptop, equipped with new software, and uses the TENZOPLUS sphygmomanometer to measure arterial blood pressure. (PNEVMOKARD (Pneumocard) is an attempt to obtain new scientific information to refine the understanding about the mechanisms used by the cardiorespiratory system and the whole body organism to spaceflight conditions. By recording (on PCMCIA cards) the crewmember's electrocardiogram, impedance cardiogram, low-frequency phonocardiogram (seismocardiogram), pneumotachogram (using nose temperature sensors), and finger photoplethismogram, the experiment supports integrated studies of (1) the cardiovascular system and its adaptation mechanisms in various phases of a long-duration mission, (2) the synchronization of heart activity and breathing factors, as well as the cardiorespiratory system control processes based on the variability rate of physiological parameters, and (3) the interconnection between the cardiorespiratory system during a long-duration mission and the tolerance of orthostatic & physical activities at the beginning of readaptation for predicting possible reactions of the crewmembers organism during the their return to ground.)

Starting a new round of RS ventilation system maintenance, the FE-1 will work in the SM, cleaning 'Group A' fan grilles.

Afterwards, Yuri will also conduct the periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various Russian segment (RS) hatchways, including the FGB-to-Soyuz tunnel, and the FGB-to-Node passageway. (This is especially important when the ventilation/circulation system has to cope with a larger crew on board, currently ten persons, and one of the two Russian SKV air conditioners off (SKV-1).)

Also later today, Behnken and Reisman will conduct a brake run-in test on the newly installed SPDM arms and then stow the arms in pre-EVA-3 configuration. (There are no temperature concerns with SPDM powered off.)

At ~4:45am tomorrow morning, Rick Linnehan (EV1) and Bob Behnken (EV2) will begin their 'campout' in the A/L with hatch closure and depressurization of the CL (Crewlock) from 14.7 to 10.2 psi, followed by mask prebreathe. Sleep time for the ISS crew begins at 6:00am. (For the Campout, fresh METOX (Metal Oxide) canisters will be installed in the A/L for CO2 control. EVA-3 will begin at ~7:25am and last approximately 6.5 hrs, ending at ~2:00am. Its major objectives are: (1) Install OTP/THA (ORU Tool Platform/Tool Holder Assembly), (2) clean up SLP, (3) transfer MISSE-6 & LWAPA (Light-Weight Adaptor Plate Assembly) and install on Columbus, (4) transfer one spare SSRMS yaw joint and two DCSUs ( (Direct Current Switching Units) from Shuttle PLB (Payload Bay) to ESP-2 (External Stowage Platform 2), (5) install two CLPAs (Camera, Light, PTU Assemblies) on SPDM, and (6) clean up worksite & ingress.)

Despite their busy timelines, all ISS crewmembers had time scheduled yesterday (as well as tonight after wakeup) for conducting their regular physical exercise.

Transfers: Middeck transfers are going well; no concerns. N2 (nitrogen) transfer is complete (~23 lbs transferred).

ATV 'Jules Verne' Status Update: On 3/14 (Friday), ATV1 performed two maneuvers. Immediately prior to the start of the second burn, PDE-4 (Propulsion Drive Electronics #4) was disabled by an FDIR (Fault Detection, Isolation & Recovery) signal reacting to a temperature threshold violation in a temperature sensor (thermistor T406). Alarms were also observed during the first burn which did not reach the FDIR threshold. Between maneuvers, ATVCC/Toulouse attempted unsuccessfully to reset the alarm counter; thus, the FDIR was triggered at the start of the second burn, disabling PDE-4. The maneuver was completed nominally on three PDEs. To date, all three incidents involving PDE disabling have been associated with other equipment FDIRs (a helium regulator and two thermistors). All three incidents have been resolved with no loss of vehicle functionality or redundancy. All ATV1 systems are performing nominally. The next burn is planned for 3/18 (Tuesday) at 6:41pm EDT. ATV1 enters parking position on 3/19 at 9:00am.

ISS Crew Sleep Shift Planning: To synchronize the ISS crew's timeline with STS-123/1J/A docking and subsequent docked activities, Peggy's, Yuri's and Leo's wake/sleep cycle is undergoing a number of shifts which started on 3/11. For the next six days, the wake/sleep shift schedule is as follows (all times EDT):

FD7 Wake: 3:30pm (3/16) - 6:00am (3/17)
Sleep: 6:00am - 2:30pm (3/17)

Wake: 3:30pm (3/17) - 6:00am (3/18)
Sleep: 6:00am - 2:30pm (3/18)

Wake: 2:30pm (3/18) - 5:00am (3/19)
Sleep: 5:00am - 1:30pm (3/19)

Wake: 1:30pm (3/19) - 5:00am (3/20)
Sleep: 5:00am - 1:30pm (3/20)

Wake: 1:30pm (3/20) - 5:00am (3/21)
Sleep: 5:00am - 1:30pm (3/21)

Wake: 1:30pm (3/21) - 4:00am (3/22)
Sleep: 4:00am - 12:30pm (3/22)

Weekly Science Update (Expedition Sixteen -- Week 21)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Radiation measurements continue to be performed in the PIRS module. Radiation measurements continue to be performed in the PIRS module. Next memory card exchange took place at 3/14.

ANITA: Completed.

BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): In progress.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Data collected during the first in-flight session have been down-linked and are under analysis. Engineering assessment provided good results: the system was calibrated within the expected 1 mm accuracy. Scientific analysis is going on. Data collected during second in-flight session are still to be downlinked. The scientific protocols were fully executed in both sessions

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): In progress.

EuTEF (European Technology Exposure Facility): DEBIE-2 showed regular link errors and was switched off on 3/9. DOSTEL: On-going science acquisition. EuTEMP: Currently inactive as planned. EVC: first EVC pictures were received on 3/6. Further commissioning was performed on 3/14 with a first successful picture acquisition run. During a second sequence with different parameters, however, again no HRD (high rate data) could be received. Under further investigation. EXPOSE: On-going science acquisition. FIPEX: Sensor units RAM3, RAM4 and ZENITH8 were switched on. However, FIPEX showed link errors and was switched off again. MEDET: Successful MEDET commanding from User Home Base on 3/10. On-going science acquisition. PLEGPAY: Further commissioning to be planned. TRIBOLAB: In Stand-by mode.

Fluid Science Laboratory (FSL): The FSL Facility commissioning is on hold pending restoration of a LAN (Local Area Network) cable connection.

GEOFLOW: Troubleshooting on FSL LAN cable repair was successfully performed. However, damage on the MIL-STANDARD bus connector was observed (two pins are missing). Start of GEOFLOW is pending further FSL troubleshooting. FSL FCE (Facility Core Element) locking was performed for 1J/A docking.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): In progress.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.

MULTIGEN-1: Completed.

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.

NUTRITION/REPOSITORY: 'Leo, thanks for your attention to detail in completing the FD30 Nutr/Rep session! The barcode information conveyed via crew notes was very helpful.'

PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): 'Peggy, thanks for completing your Actiwatch download. We have downlinked your data and sent it to the PI. We have placed on your task list your last scheduled sleep logging session for next week. Thanks for the extra sleep logging.'

SOLAR (Solar Monitoring Observatory): Anomaly on platform pointing mode under further investigation. No science acquisition possible so far.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Thank you, Peggy, for completing the additional SWAB session off the Voluntary Science list. The extra data collected will assist the PI in a comparison between the previous samples in the USL and those you recently collected from the new modules that they would not have originally been able to obtain. '

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

ULTRASOUND: In progress.

WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Further attempts made to reduce condensation in ECs (Experiment Containers) by continuous flushing and by flushing only 2 EC's at a time (to increase the flow / pressure) did not reduce the condensation significantly. The condensation presented a major challenge on high-resolution picture quality for the crew activity on 3/10. To recover, an additional video session was proposed during the fixation because the fixation fluid clears the EC windows. During the ground commanding after the photo activity, two BIOLAB (BLB) anomalies occurred which lead to a loss of centrifugation and temperature control for ~4:20 versus the 2 hours that was originally foreseen. Fixation & washing was planned during the night from 3/11 to 3/12. Multiple BLB anomalies occurred on both rotors. Troubleshooting did not allow to perform any fixation on 3/12 or 3/13. Rotor B was blocked and could not provide a 1-g stimulus since 3/11. The 4 EC's from Rotor A (0-g condition) were transferred to BLB TCU at +4degC to slow down the growth and will be downloaded on 1J/A. 4 EC's of Rotor B have yet to be retrieved by the crew from the blocked Rotor B. Anomalies during WAICO-Run#1 imply major impact on science return.

CEO (Crew Earth Observation): Ongoing.

No CEO photo targets uplinked for today.

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use